A copper rod and a steel rod are heated. At 0 ◦C the copper rod has a length of LC, the steel one has a length of LS. When the rods are being heated or cooled, a difference of 3.84 cm is maintained between their lengths. If αC = 1.7 × 10−5 ( ◦C)−1 and αS = 1.2 × 10−5 ( ◦C)−1 , find the value of LC. Answer in units of cm.
Answer #1

Since given that a difference of 3.84 cm is maintain between their lengths, So change in length of both rods with temperature will remain same

dLc = dLs

dLc = Lc*alphaC*dT

dLs = Ls*alphaS*dT

Which gives

Lc*alphaC*dT = Ls*alphaS*dT

Lc*alphaC = Ls*alphaS

Lc = Ls*(alphaS/alphaC)

Now given that

Ls - Lc = 3.84 cm = 0.0384

(See that we will use Ls - Lc, because steel's temperature coefficient is lower than that of copper, So length of steel will be higher than that copper)

Ls = Lc + 0.0384

Using above value

Lc = Ls*(alphaS/alphaC)

Lc = (Lc + 0.0384)*(alphaS/alphaC)

Lc*(1 - alphaS/alphaC) = 0.0384*(alphaS/alphaC)

Lc*(1 - 1.2*10^-5/(1.7*10^-5)) = 0.0384*1.2*10^-5/(1.7*10^-5)

Lc*0.294 = 0.0271

Lc = 0.0271/0.294

Lc = 0.0922 m = 9.22 cm

Please Upvote.